OpenStudy (peterg2001):

A pentagon has interior angles that sum to _____. 360° 540° 720° 900° @Molly_Qian

OpenStudy (anonymous):

Each angle is 108 degrees. Times that by five and they ye go. 540

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (peterg2001):

not 540 for sure i tried that

OpenStudy (anonymous):

To add to what @Mech123 said, he came to that 108 degrees by knowing that the exterior angles all add up to 360, so we know each angle is 360/5=72. Then the interior angle is 180-72=108. There's 5 of them so 5*108=540.

OpenStudy (anonymous):

Yea what he said

OpenStudy (peterg2001):

sorry u were right thanks all